OpenStudy (anonymous):

Use the Ratio Test to find the interval of convergence of SUM [(-3x)^n/3n+1 ]. I have -1/3

OpenStudy (zarkon):

your interval is a little off

OpenStudy (anonymous):

What did you get??

OpenStudy (zarkon):

\[\left(-\frac{1}{3},\frac{1}{3}\right]\]

OpenStudy (zarkon):

\[-1/3<x\le 1/3\]

OpenStudy (anonymous):

Ok ,s o how do you know that 1/3 is included??

OpenStudy (zarkon):

alternating series test

OpenStudy (anonymous):

Wellllll........i did the AST to get those two answers. But should I plug those into the equation?

OpenStudy (zarkon):

you need to plug -1/3 and 1/3 into the orignial sum for x and then test for convergence

OpenStudy (anonymous):

oooooooooooh, im sorry. I understand. What I did was called the Ratio test. But now I need to use the AST.

OpenStudy (zarkon):

you need to check the end points...with one of them (1/3) you can use the ast